OpenStudy (anonymous):

PLEASE HELP ME If a roof has a pitch of 5 to 11, how high will the roof rise over a 33-foot run?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve for x 5/11 = x/33

OpenStudy (wolf1728):

If the pitch is 5 to 11 that means the roof rises 5 feet for every 11 horizontal feet.

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

????

OpenStudy (wolf1728):

The answer is not 4

jimthompson5910 (jim_thompson5910):

5/11 = x/33 5/11 = 4/33 ... plug in x = 4 is that equation true?

OpenStudy (anonymous):

LOL i mean 15

jimthompson5910 (jim_thompson5910):

x = 15 is correct